User name S. Sirca
Log entry time 00:00:21 on October28,2001
Entry number 75248
keyword=Right HRS floor-marks
When the right spectrometer was moved from 56.217 deg to 59.97 deg,
the pointer on the additional "ruler-gauge" on the TV screen moved from
100.5 to 107.4.